# Bounds on decaying sterile neutrinos via magnetic dipole moment from COB intensity

### v2

@@ IV Results and Discussions
In the previous section, we discussed that the intensity of photon production by the sterile neutrino radiative decays which is linked to the neutrino magnetic moment. Here, we investigate the reported COB anomalous intensities, that is $8.06\pm 1.92\,\rm nW/m^{2}/sr$ Lauer et al. (2022) and $2.99\pm 2.03\leavevmode\nobreak\ \,\rm nW/m^{2}/sr$ Postman et al. (2024), and analyze them within the framework of the neutrino magnetic moment to derive upper bounds on the mass and magnetic moment. From hereinafter we will represent the reported intensities $8.06\pm 1.92\,\rm nW/m^{2}/sr$ and $2.99\pm 2.03\leavevmode\nobreak\ \,\rm nW/m^{2}/sr$ as $I_{806}$ Lauer et al. (2022) and $I_{299}$ respectively Postman et al. (2024), unless stated otherwise. Specifically, we investigate three possible scenarios: sterile-sterile, sterile-active, and active-active neutrino magnetic moments.
@@ IV.1 Sterile to sterile neutrino decay
Figure 2: Top panel: The minimum required decay width to obtain $I_{806}$ and $I_{299}$ shown in red dashed and blue solid lines, respectively, for masses $1\leavevmode\nobreak\ \rm keV-40\leavevmode\nobreak\ \rm keV$. Bottom panel: The corresponding upper bounds on $d_{\rm NN\gamma}$ for the anomalous intensities $I_{806}$ Lauer et al. (2022) and $I_{299}$ Postman et al. (2024).
@@
Figure 3: This figure illustrates the upper bounds on $d_{\rm NN\gamma}$ versus the mass of the sterile neutrino $(m_{1})$. The red dashed and blue solid line denote $d_{NN\gamma}$ obtained for intensity $I_{806}$ Lauer et al. (2022) and $I_{299}$ Postman et al. (2024). The lavender-shaded region represents the excluded values of $d_{\rm NN\gamma}-m_{1}$ values from X-ray observations Natwariya and Nayak (2022).
@@
In this section, we calculate the required decay rate and the transition magnetic moment for the anomalous intensities reported by LORRI Lauer et al. (2022); Postman et al. (2024). We consider the keV mass range sterile neutrino as a dark matter candidate, which decays into another sterile neutrino and photon through the radiative decay process. We note that the sterile neutrinos with mass $<0.4\,\rm keV$ are excluded as a dark matter candidate by a conservative application of Tremaine-Gunn bound Abazajian and Koushiappas (2006). Now, to calculate $I_{\lambda}$ required to explain the anomalous intensity of COB, we need to determine the energy of photons $(\Delta m)$, as shown in Eq. (8). The LORRI observation reported flux of photons at pivot wavelength $0.608\,\mu m\,(\sim 2\,\rm eV)$ Lauer et al. (2022). Therefore, $\Delta m$ should be in the range $2\,\rm eV-10\,\rm eV$, such that LORRI’s band could observe these redshifting photons today. We restrict the energies of photons in this range, as the photons produced at $z\gtrsim 4$ will get absorbed by the intergalactic gas medium Madau (1995); Inoue et al. (2014).
@@
To illustrate the variation of $I_{\lambda}$ with $\lambda_{\rm obs}$, we considered $1\,\rm keV$ and $3\,\rm keV$ sterile neutrinos decaying at $10^{-22}\,\rm sec^{-1}$ and $10^{-21}\,\rm sec^{-1}$, respectively— shown in blue solid lines and red dashed lines, respectively, in Fig. 1. In this figure, the cutoff vertical line for wavelength $\lambda_{e}$ is the wavelength of photons observed today. The green-shaded region represents the shape responsivity of LORRI’s camera with respect to the observed wavelength Weaver et al. (2020); Bernal et al. (2022); Porras-Bedmar et al. (2024). Further, we fixed the intensities to the LORRI’s reported values, shown in Fig. 1. We first fix the obtained intensity to $I_{806}$ Lauer et al. (2022), for $1-20\,\rm keV$ sterile neutrinos decaying in the range $\Gamma_{m_{1}}\sim\left(10^{-21}-10^{-20}\right)\,\rm sec^{-1}$, as depicted in the red dashed lines. Similarly, the blue solid lines show $1-40\,\rm keV$ sterile neutrinos decaying at $\Gamma_{m_{1}}\sim\left(10^{-22}-10^{-21}\right)\,\rm sec^{-1}$ producing the LORRI’s reported intensity $I_{299}$ Postman et al. (2024). We note that the intensity curves at lower $\lambda_{\rm obs}$ values suggests production photons with energy $\Delta m$ for the aforementioned $m_{1}$ and $\Gamma_{m_{1}}$ originating from higher redshifts. From Eq. (8), we can observe that high energetic photons originating at higher redshifts require smaller $\Gamma_{m_{1}}$. Therefore, a sterile neutrino with a large decay width is required to produce less energetic photons in order to contribute to the anomalous COB intensities. In the next section, we present the minimum required $\Gamma_{m_{1}}$, and hence the $d_{\rm NN\gamma}$, to obtain the reported COB anomalous intensities for different $m_{1}$ values.
@@
In the top panel of Fig. (2), we plot the required minimum decay width for different masses to obtain the anomalous intensities. The red dashed and blue solid lines represent the minimum decay width for masses $1-40\,\rm keV$ in order to obtain the reported intensities in $I_{806}$ and $I_{299}$, respectively. We find that sterile neutrinos of mass $3\,\rm keV$ and $1\,\rm keV$ require $1\times 10^{-21}\,\rm sec^{-1}$ and $8.54\times 10^{-22}\,\rm sec^{-1}$ for intensities reported in $I_{8.06\pm 1.92}$ and $I_{2.99\pm 2.03}$, respectively. The decay width of the sterile neutrino increases as we increase the mass. This can be analysed from the fact that $I_{\lambda}\propto\Gamma_{m_{1}}/m_{1}$— shown in Eq. (8). We restrict our analysis to $1\leq m_{1}/\rm keV\leq 40$ because a further increase in $m_{1}$ requires an increase in $\Gamma_{m_{1}}$. For instance, the required $\Gamma_{m_{1}}$ for $m_{1}\gtrsim 40\,\rm keV$ requires $\gtrsim 10^{-20}\,\rm sec^{-1}$ for $I_{806}$, which become comparable to $\sim t^{-1}_{\rm U}$ which requires a consideration of variation in the dark matter density— as mentioned in section (III). We then plot the minimum values of $d_{\rm NN\gamma}$, in the bottom panel, for the above obtained $\Gamma_{m_{1}}$ values — shown in the red dashed and blue solid lines. We find that under the conditions $m_{1}+m_{2}\sim 2m_{1}$ and $m_{1}-m_{2}=\Delta m$, the decay width Eq. 2 becomes $16|d_{\rm NN\gamma}|^{2}\,\Delta m^{3}/\pi$. As mentioned earlier, $\Delta m$ takes values of $2\,\rm eV-10\,\rm eV$ for the observed flux of photons. Therefore, to calculate upper bounds on $d_{\rm NN\gamma}$, we consider the minimum values of $\Gamma_{m_{1}}$ for the corresponding $m_{1}$ values while fixing $\Delta m=10\,\rm eV$. We find that the transition magnetic moment $d_{\rm NN\gamma}/\rm eV^{-1}$ takes values of $(3\times 10^{-11}-10^{-8})\mu_{B}$ and $(3\times 10^{-12}-10^{-9})\mu_{B}$ for mass $1\,\rm keV-40\,\rm keV$— as shown in the red dashed and blue solid lines, respectively. Further in Fig. (3), we present the upper bounds on $d_{\rm NN\gamma}$ in the presence of existing X-ray bounds on $\Gamma_{m_{1}}-m_{1}$ values as the lavender coloured region Natwariya and Nayak (2022). In the next section, we derive constraints on the sterile-to-active transition magnetic dipole moment $(d_{\nu N\gamma})$ by considering the COB anomalous intensities $I_{806}$ and $I_{299}$.
@@ IV.2 Sterile to active neutrino decay
In this section, we consider the $\mathcal{O}$(eV) mass scale of sterile neutrino, which is decaying into active neutrino and photon via active-to-sterile transition magnetic dipole moment. This radiative decay process can be expressed as $N_{1}\to\nu_{a}+\gamma$. The wavelength of the emitted photon $(\lambda_{e})$ can be calculated from Eq. (4). The term $\lambda_{e}$ depends on both the mass of sterile and active neutrino. Here, we fix the active neutrino mass to be $0.12\,\rm eV$. As explained earlier in section (IV.1), the required energy of the photons should be between $2\,\rm eV-10\,\rm eV$ such that these redshifting photons will be observed today as an excess radiation in the COB. Using these conditions, we find that the mass of the sterile neutrino will be in the range of $4\,\rm eV-20\,\rm eV$. The upper bound on the mass of the sterile neutrino to be a dark matter candidate, derived from the Traemaine-Gunn bound, is $0.4$ keV Abazajian and Koushiappas (2006). Therefore, we examine the $\mathcal{O}$(eV) scale decaying sterile neutrinos, which may be considered as dark radiation. In the standard scenario case, only three neutrino species contribute to the $\rm N_{eff}^{SM}$ which is $3.044$. However, in a nonstandard scenario, the effective neutrino species can be written as $\rm N_{eff}=N_{eff}^{SM}+\Delta N_{eff}$, where $\Delta\rm N_{eff}$ comes from the additional relativistic components that take part in the dark radiation Abazajian and Koushiappas (2006). Dark radiation measures the amount of radiation energy contributed by the relativistic species except photons. Usually, neutrino oscillation anomalies give the idea for searching the new additional relativistic particles in the cosmos Abazajian and Koushiappas (2006). Motivated by the neutrino oscillation experiment for explaining the dark radiation, the mass range of sterile neutrino should be $m_{s}\lesssim\rm eV$ Drewes (2013). So here, we take the mass range $(4\,\rm eV-20\,\rm eV)$ of decaying sterile neutrino, which can be regarded as dark radiation.
Figure 4: (a): Represent the variations in specific intensity $I_{\lambda}$ with respect to $\lambda_{\rm obs}$, for $12$ eV sterile neutrinos decaying at $10^{-23}\,\text{sec}^{-1}$ and $10^{-22}\,\text{sec}^{-1}$, shown in blue solid and red dashed lines, respectively. (b): Shows variations in $I_{\lambda}$ for $12$ eV sterile neutrinos of masses with decay rates $10^{-23}-10^{-22}\,\rm sec^{-1}$ for fixed intensities $I_{806}$ and $I_{299}$, as shown in the red dashed and blue solid lines, respectively. The green-shaded area illustrates the responsivity curve of LORRI for each detected wavelength.
The decay width for a decaying sterile neutrino via sterile-to-active transition magnetic moment is given by Beltrán et al. (2024)
$\Gamma_{m_{1}}(m_{1}\rightarrow m_{2}+\gamma)=\frac{|d_{\rm\nu N\gamma}|^{2}}{2\pi}m_{1}^{3}\,,$ (9)
where $d_{\nu\rm N\gamma}$, $m_{1}$, and $m_{2}$ represent the active-to-sterile transition magnetic dipole moment and the mass of the sterile and active neutrino, respectively. On the contrary to the decay width shown in Eq. (2), here $\Gamma_{m_{1}}$ is independent of the mass difference between the sterile and active neutrino. We then write the expression for the mean specific intensity $I_{\lambda}$ of photons produced from eV-scale decaying sterile neutrino as
$I_{\lambda}=x\,\frac{c}{4\pi}\frac{{\Omega_{s}}{\rho_{c}}{c^{2}}\Gamma_{m_{1}}}{\lambda_{obs}(1+z)H(z)}.$ (10)
Here, the term $x=1/2\,\left(1-m_{2}^{2}/m_{1}^{2}\right)$ is similar to the one shown in Eq. (4). Whereas, $\Omega_{s}$ represents the density parameter of sterile neutrinos, which can be expressed as $\Omega_{s}h^{2}=m_{1}/94.1\,\rm eV$ for a nonthermal scenario, which comes from the “Dodelson-Widrow Mechanism” Dodelson and Widrow (1994).
Using Eq. (10), we calculate $I_{\lambda}$ for $(4\,\rm eV-20\,\rm eV)$ mass range of sterile neutrinos, that can contribute to the reported intensities by LORRI in the wavelength range of $(0.4\,\rm\mu m-0.9\,\rm\mu m)$. For a fixed $m_{1}=12\,\rm eV$, we considered decay widths $\Gamma_{m_{1}}$ of $10^{-23}\,\rm sec^{-1}$ and $-10^{-22}\,\rm sec^{-1}$ to plot variations in $I_{\lambda}$ for different $\lambda_{\rm obs}$ values— shown in blue solid and red dashed lines, respectively, in Fig. 4. Further in Fig. 4, we vary $m_{1}$ in the range $(4\,\rm eV-20\,\rm eV)$ and $\Gamma_{m_{1}}$ within $10^{-23}-10^{-22}\,\rm sec^{-1}$, while fixing the total intensity of photons to $I_{806}$ and $I_{299}$— as shown in red dashed and blue solid lines, respectively. Below, we calculate the decay rate for mass $m_{1}$ required to obtain both the COB intensities and derive upper bounds on sterile-to-sterile transition magnetic dipole moment.
Figure 5: Top panel: The minimum required decay rate to put constraint on sterile-to-active transition magnetic dipole moment ($d_{\nu N\gamma}$) by taking into consideration the contribution of the anomalous intensity of COB $2.99\pm 2.03\,\rm nW/m^{2}/sr$ Postman et al. (2024) and the anomalous COB excess intensity $8.06\pm 1.92\,\rm nW/m^{2}/sr$ Lauer et al. (2022) reported by LORRI for sterile neutrinos with masses ($4\leavevmode\nobreak\ \rm eV-20\leavevmode\nobreak\ \rm eV)$. Bottom panel: The corresponding minimum magnetic coupling strength $(d_{\rm\nu N\gamma})$ in $\text{eV}^{-1}$.
In the top panel of Fig. 5, the blue solid and red dashed lines show the minimum decay width $\Gamma_{m_{1}}$ for sterile neutrino $4\leavevmode\nobreak\ \text{eV}\leq m_{1}\leq 20\leavevmode\nobreak\ \text{eV}$ to obtain the total intensities of $I_{299}$ and $I_{806}$, respectively. We find that $\Gamma_{m_{1}}$ increases for heavier sterile neutrinos. This can be analyzed from the fact that the specific intensity $(I_{\lambda})$ depends directly on the decay width while mass enters only through redshift $z(\lambda_{\rm obs},m_{1})$ which can only alter the Hubble parameter $H(z)$— as shown in Eq. (10). Thus, larger mass increases $z(\lambda_{\rm obs},m_{1})$, suggesting an early decay of sterile neutrino. However, as the intensity of photons redshifts more for an early decay, a larger decay width is required to obtain the required intensity today. Further, the solid blue and dashed red lines in the bottom panel show upper bounds on active-to-sterile transition magnetic moment $d_{\nu N\gamma}$ derived for the above blue solid line and red dashed line respectively using Eq. (9). We observe that $d_{\nu N\gamma}$ decreases with the increase of the mass of the decaying sterile neutrino. This can be analysed from Eq. (9), where $d_{\nu N\gamma}$ is inversely proportional to the mass of decaying sterile neutrinos. We find that, the required $d_{\nu N\gamma}$ to explain the anomalous intensities of $I_{806}$ and $I_{299}$ lies between $3.33\times 10^{-12}\,\rm\mu_{B}-10^{-11}\,\rm\mu_{B}$ and $3.33\times 10^{-13}\,\rm\mu_{B}-10^{-12}\,\rm\mu_{B}$, respectively, for sterile neutrino mass $4\,\rm eV\leq m_{1}\leq 20\,\rm eV$.
Currently, the upper bound on the active-to-sterile transition dipole moment $d_{\rm\nu N\gamma}$ for sterile neutrino’s mass $4\,\rm eV\leq m_{1}\leq 20\,\rm eV$ remains unconstrained. There are several experimental upper bounds on the active-to-sterile transition magnetic dipole moment. Recent studies of XENON1T found that the upper bound on $d_{\nu N\gamma}\lesssim 6\times 10^{-11}\mu_{B}$ for several MeV mass range of sterile neutrino Shoemaker et al. (2021). However, the recent analysis of global clusters sets an upper limit as $d_{\nu N\gamma}\lesssim 1.2\times 10^{-12}\mu_{B}$ Cañas et al. (2016).
IV.3 Active-to-active neutrino magnetic moment

### v3

@@ IV Results and Discussions
In the previous section, we discussed that the intensity of photon production by the sterile neutrino radiative decays which is linked to the neutrino magnetic moment. Here, we investigate the reported COB anomalous intensities, that is $8.06\pm 1.92\,\rm nW/m^{2}/sr$ Lauer et al. (2022) and $2.99\pm 2.03\leavevmode\nobreak\ \,\rm nW/m^{2}/sr$ Postman et al. (2024), and analyse them within the framework of the neutrino magnetic moment to derive an upper bound on the transition magnetic moment. From hereinafter we will represent the reported intensities $8.06\pm 1.92\,\rm nW/m^{2}/sr$ and $2.99\pm 2.03\leavevmode\nobreak\ \,\rm nW/m^{2}/sr$ as $I_{806}$ Lauer et al. (2022) and $I_{299}$ Postman et al. (2024), unless stated otherwise. Specifically, we investigate the possible scenario: the sterile-to-sterile transition magnetic moment.
@@ IV.1 Sterile to sterile neutrino decay
Figure 2: Top panel: The minimum required decay width to obtain $I_{806}$ and $I_{299}$ shown in red dashed and blue solid lines, respectively, for masses $1\leavevmode\nobreak\ \rm keV-40\leavevmode\nobreak\ \rm keV$. Bottom panel: The corresponding upper bounds on $d_{\rm NN\gamma}$ for the anomalous intensities $I_{806}$ Lauer et al. (2022) and $I_{299}$ Postman et al. (2024). The colour band represents the change of the sterile-to-sterile transition magnetic moment $(d_{NN\gamma})$ in $\rm eV^{-1}$ versus the mass of sterile neutrino $(m_{1})$ in keV scale, where, $\Delta m$ vary from $2-10\,\rm eV$. The green one represents the variation of $d_{NN\gamma}$ due to the anomalous rest intensity of COB, $2.99\pm 2.03\,\rm nW/m^{2}/sr$. Whereas, the yellow one represents the same variation of $d_{NN\gamma}$ due to the anomalous COB excess intensity $8.06\pm 1.92\,\rm nW/m^{2}/sr$.
@@
Figure 3: The yellow and green colour bands illustrate the variation of $d_{\rm NN\gamma}$ versus the mass of the sterile neutrino $(m_{1})$ with the changes of $\Delta m$, which vary from $2\,\rm eV$ to $10\,\rm eV$ . These two-colour bands denote the changes of $d_{NN\gamma}$ obtained from anomalous intensities of COB, $I_{806}$ Lauer et al. (2022) and $I_{299}$ Postman et al. (2024), respectively. The cyan-shaded region represents the excluded values of $d_{\rm NN\gamma}-m_{1}$ values from X-ray observations Natwariya and Nayak (2022).
@@
In this section, we calculate the required decay width and the transition magnetic moment for the anomalous intensities reported by LORRI Lauer et al. (2022); Postman et al. (2024). We consider the keV mass range sterile neutrino as a dark matter candidate, which decays into another sterile neutrino and photon through the radiative decay process. We note that the sterile neutrinos with mass $0.4\,\rm keV$ are excluded as a dark matter candidate by a conservative application of the Tremaine-Gunn bound Abazajian and Koushiappas (2006). Now, to calculate $I_{\lambda}$ required to explain the anomalous intensity of COB, we need to determine the energy of photons $(\Delta m)$, as shown in Eq. (8). The LORRI observation reported flux of photons at pivot wavelength $0.608\,\mu m\,(\sim 2\,\rm eV)$ Lauer et al. (2022). Therefore, $\Delta m$ should be in the range $2\,\rm eV-10\,\rm eV$, such that LORRI’s band could observe these redshifting photons today. We restrict the energies of photons in this range, as the photons produced at $z\gtrsim 4$ will get absorbed by the intergalactic gas medium Madau (1995); Inoue et al. (2014).
@@
To illustrate the variation of $I_{\lambda}$ with $\lambda_{\rm obs}$, here we consider $1\,\rm keV$ and $3\,\rm keV$ masses of sterile neutrinos decaying at $10^{-22}\,\rm sec^{-1}$ and $10^{-21}\,\rm sec^{-1}$, respectively— shown in blue solid lines and red dashed lines, respectively, in Fig. 1. In this figure, the cutoff vertical line for wavelength $\lambda_{e}$ is the wavelength of photons observed today. The green-shaded region represents the shape responsivity of LORRI’s camera with respect to the observed wavelength Weaver et al. (2020); Bernal et al. (2022); Porras-Bedmar et al. (2024). Furthermore, we adjust the intensities to match LORRI’s reported values, as shown in Fig. 1. We first fix the obtained intensity to $I_{806}$ Lauer et al. (2022), for $1-20\,\rm keV$ sterile neutrinos decaying in the range $\Gamma_{m_{1}}\sim\left(10^{-21}-10^{-20}\right)\,\rm sec^{-1}$, as depicted in the red dashed lines. Similarly, the blue solid lines show $1-40\,\rm keV$ sterile neutrinos decaying at $\Gamma_{m_{1}}\sim\left(10^{-22}-10^{-21}\right)\,\rm sec^{-1}$ producing the LORRI’s reported intensity $I_{299}$ Postman et al. (2024). We note that the intensity curves at lower $\lambda_{\rm obs}$ values suggest the production of photons with energy $\Delta m$ for the aforementioned $m_{1}$ and $\Gamma_{m_{1}}$ originating from higher redshifts. From Eq. (8), we can observe that high energetic photons originating at higher redshifts require smaller $\Gamma_{m_{1}}$. Therefore, a sterile neutrino with a large decay width is required to produce less energetic photons in order to contribute to the anomalous COB intensities. In the next section, we present the minimum required $\Gamma_{m_{1}}$, and hence the $d_{\rm NN\gamma}$, to obtain the reported COB anomalous intensities for different $m_{1}$ values.
@@
In the top panel of Fig. (2), we plot the required minimum decay width for the mass range of sterile neutrinos $1\,\rm keV-40\,\rm keV$ to obtain the anomalous intensities of COB. The red dashed and blue solid lines represent the minimum decay width for masses $1-40\,\rm keV$ to obtain the reported intensities in $I_{806}$ and $I_{299}$, respectively. We find that sterile neutrinos of mass $3\,\rm keV$ and $1\,\rm keV$ require $1\times 10^{-21}\,\rm sec^{-1}$ and $8.54\times 10^{-22}\,\rm sec^{-1}$ for intensities reported in $I_{8.06\pm 1.92}$ and $I_{2.99\pm 2.03}$, respectively. The decay width of the sterile neutrino increases as we increase the mass. This can be analysed from the fact that $I_{\lambda}\propto\Gamma_{m_{1}}/m_{1}$— shown in Eq. (8). We restrict our analysis to $1\leq m_{1}/\rm keV\leq 40$ because a further increase in $m_{1}$ requires an increase in $\Gamma_{m_{1}}$. For instance, the required minimum decay width $\Gamma_{m_{1}}$ for $m_{1}\gtrsim 40\,\rm keV$ requires $\gtrsim 10^{-20}\,\rm sec^{-1}$ for $I_{806}$, which becomes comparable to $\sim t^{-1}_{\rm U}$ which requires a consideration of variation in the dark matter density— as mentioned in section (III). We then plot the minimum values of $d_{\rm NN\gamma}$, in the bottom panel, for the above obtained $\Gamma_{m_{1}}$ values — shown in the red dashed and blue solid lines. We find that under the conditions $m_{1}+m_{2}\sim 2m_{1}$ and $m_{1}-m_{2}=\Delta m$, the decay width Eq. 2 becomes $16|d_{\rm NN\gamma}|^{2}\,\Delta m^{3}/\pi$. As mentioned earlier, $\Delta m$ takes values of $2\,\rm eV-10\,\rm eV$ for the observed flux of photons. We find that the sterile-to-sterile transition magnetic moment $d_{\rm NN\gamma}/\rm eV^{-1}$ takes values of $(3\times 10^{-12}-10^{-8})$ and $(3\times 10^{-13}-10^{-9})$ for mass $1\,\rm keV-40\,\rm keV$— as shown in the red dashed and blue solid lines, respectively. In Fig. 2, we illustrate the colour band of $\Delta m$ variation from $2\,\rm eV$ to $10\,\rm eV$ for the two anomalous intensities of COB. This colour band represents the changes of the sterile-to-sterile transition magnetic moment $(d_{NN\gamma})$ in $\rm eV^{-1}$ versus the mass of the sterile neutrino $(m_{1})$ for the changes of the minimum required decay width of the sterile neutrino. The green colour band represents the changes of the sterile-to-sterile transition magnetic moment $(d_{NN\gamma})$ for the anomalous rest intensity of COB, $2.99\pm 2.03\,\rm nW/m^{2}/sr$. Whereas, the yellow colour band region represents the changes $d_{NN\gamma}$ for the anomalous COB excess intensity $8.06\pm 1.92\,\rm nW/m^{2}/sr$. From the decay width Eq. (3), we can see that the sterile-to-sterile transition magnetic moment $(d_{NN\gamma})$ depends directly on the decay width of the sterile neutrino. We observe that for the smaller mass difference, $\Delta m=2\leavevmode\nobreak\ \,\rm eV$, the sterile-to-sterile transition magnetic moment $(d_{NN\gamma})$ approaches up to $10^{-13}\,\rm eV^{-1}$ and the minimum required decay width $(\Gamma_{m_{1}})$ reaches up to $10^{-21}\,\rm sec^{-1}$ for the anomalous COB excess intensity $8.06\pm 1.92\,\rm nW/m^{2}/sr$ – as shown in the red dashed line. Whereas, for the larger mass difference between two heavy sterile states, $\Delta m=10\leavevmode\nobreak\ \,\rm eV$, the minimum required decay width of the sterile neutrino $(\Gamma_{m_{1}})$ increases and reaches up to $\sim 10^{-19}\,\rm sec^{-1}$ for the mass range of the sterile neutrino $m_{1}\geq 40\,\rm keV$. Similarly, the sterile-to-sterile transition magnetic moment $(d_{NN\gamma})$ also increases and approaches up to $10^{-8}\,\rm eV^{-1}$. So, we observe that for the mass range of the sterile neutrino $m_{1}\geq 40\,\rm keV$ or the larger mass difference between two heavy sterile states $\Delta m=10\,\rm eV$, we obtain that the minimum required decay width of the sterile neutrino approaches up to $t_{\,\rm U}\sim 10^{18}\,\rm sec$. We illustrate the changes $d_{NN\gamma}$ due to the changes of mass difference $\Delta m$ from $2\,\rm eV$ to $10\,\rm eV$ in this Fig. 2. Further in Fig. (3), we present the upper bound on the $d_{\rm NN\gamma}$ in the presence of existing X-ray bounds on $\Gamma_{m_{1}}-m_{1}$ values as the cyan-coloured region Natwariya and Nayak (2022). For instance, the upper bound on the sterile-to-sterile transition magnetic moment $d_{NN\gamma}$ from X-ray observations is approximately $10^{-6}\,\rm eV^{-1}$ for the mass of sterile neutrino $5\,\rm keV$. In contrast, for the larger value of sterile neutrino mass, $40\,\rm keV$, the upper constraint on $(d_{NN\gamma})$ yields approximately $10^{-2}\,\rm eV^{-1}$ .
@@ IV.2 Sterile to active neutrino decay
In this section, we examine light sterile neutrinos in the mass range $m_{s}\in[4,\,20]\,\mathrm{eV}$. These neutrinos behave as dark radiation rather than cold dark matter. Consequently, their abundance is typically computed using the Dodelson–Widrow (DW) mechanism, given by $\Omega_{s}h^{2}=\frac{m_{s}}{94.1\,\mathrm{eV}}$. However, for sterile neutrino masses within this range, the DW mechanism is ruled out by cosmological observations, including data from the Cosmic Microwave Background (CMB) and large-scale structure formation Xu et al. (2022).
Assuming a conservative minimum relic temperature of $T_{X}=0.91\,\mathrm{K}$, stringent cosmological constraints exist on the mass of thermal light relics. Ref. Xu et al. (2022) analyzes four categories of such relics—real scalars, Weyl fermions, Dirac fermions, and vectors. Their analysis shows that minimally coupled Weyl fermions with $m_{X}>2.3\,\mathrm{eV}$, real scalars with $m_{X}>11\,\mathrm{eV}$, and Dirac fermions with $m_{X}>1.1\,\mathrm{eV}$ are excluded at the 95% confidence level Aker et al. (2019). These bounds are notably more stringent—by factors of 2 to 5—than those reported in previous literature Xu et al. (2022).
Furthermore, results from CMB analyses, weak gravitational lensing Abbott et al. (2022), and galaxy surveys strongly constrain light thermal relics with masses $m_{X}\gtrsim 0.1\,\mathrm{eV}$. Therefore, the observed excess in the Cosmic Optical Background (COB) intensity cannot be attributed to the decay of sterile neutrinos into active neutrinos within the considered mass range of $4\,\mathrm{eV}\leq m_{s}\leq 20\,\mathrm{eV}$.
IV.3 Active-to-active neutrino transition magnetic moment